# Make Autoregressive Great Again: Diffusion-Free Graph Generation with Next-Scale Prediction

### v1

@@ Abstract
Autoregressive models are popular generative models due to their speed and properties. However, they require an explicit sequence order, which contradicts the unordered nature of graphs. In contrast, diffusion models maintain permutation invariance and enable one-shot generation but require up to thousands of denoising steps and additional features, leading to high computational costs. Inspired by recent breakthroughs in image generation—especially the success of visual autoregressive methods [1, 2]—we propose MAG, a novel diffusion-free graph generation framework based on next-scale prediction. By leveraging a hierarchy of latent representations, the model progressively generates scales of the entire graph without the need for explicit node ordering. Extensive experiments on both generic and molecular graph datasets demonstrate that MAG delivers competitive performance compared to state-of-the-art methods, achieving up to three orders of magnitude in speedup during inference.
@@ 1 Introduction
Despite significant progress in generative models for language and images, graph generation remains challenging due to its inherent combinatorial nature. Specifically, graphs are naturally high-dimensional and discrete with varying sizes, contrasting with the continuous space and fixed-size techniques that cannot be directly applied to them. Furthermore, rich substructures in graphs necessitate an expressive model capable of capturing higher-order motifs and interactions.
Several graph generative models have been proposed to address some of these challenges, undertaking approaches based on autoregressive (AR) models [3], varational autoencoders (VAE) [4], generative adversarial networks (GAN) [5], and diffusion models [6, 7, 8]. Among these, diffusion and autoregressive models stand out with superior performance and interesting properties [9, 10, 11].
On one hand, diffusion models offer the ability to achieve exchangeable probability in combination with permutation equivariant networks, under certain conditions [6, 7]. However, given the high-dimensional nature of graphs and their complex internal dependencies, directly modeling the joint distribution of all nodes and edges presents significant challenges. Recent work by Zhao et al. [10] demonstrates that, as a cost for permutation invariance, capturing the full joint distribution and solving the transformation difficulty via diffusion requires thousands of sampling and denoising steps as well as additional node-, edge-, and graph-level features, such as eigenvectors, to break symmetries and achieve high generation quality, rendering diffusion a promising but very expensive approach.
On the other hand, in addition to their improved efficiency, studies into the success of AR models have highlighted their scalability and generalizabilty. The former, as exemplified by scaling laws [12], allows the prediction of large models’ performance to better guide resource allocation during training, while the latter, as evidenced by zero-shot and few-shot learning [13], underscores the adaptability of unsupervised-trained models to unseen tasks. Furthermore, the aforementioned discrete and varying-size properties of graphs suggest a promising fit for autoregressive modeling. However, unlike natural language sentences with an inherent left-to-right ordering, the order of nodes in a graph must be explicitly defined for unidirectional autoregressive learning. Previous AR methods [3, 14, 15] attempted to enforce node ordering or approximate the marginalization over permutations such that, once flattened, one can train an autoregressive model to maximize its likelihood via next-token prediction. This gives rise to the three main issues of autoregressive graph generation:
@@
    Long-range bottleneck. Generating nodes one at a time results in sequential dependencies that make it difficult to capture long-range interactions and global structures. Furthermore, early predictions errors can propagate throughout the generation process, and dependencies may decay when working on larger graphs.
@@
    Counterintuitive node ordering. Enforcing a specific node ordering is at odds with the inherent invariance of graphs, forcing the model to learn over a specific ordering, thereby increasing complexity and reducing generalization.
@@
    Inefficiency. Generating a graph sequence $(x_{1},x_{2},...,x_{N\times N})$ with a conventional self-attention transformer incurs $\mathcal{O}(n^{2})$ autoregressive steps and $\mathcal{O}(n^{3})$ computational cost.
Drawing inspiration from recent breakthroughs in image generation—notably the Visual Autoregressive Modeling (VAR) framework [1, 2]—our proposed method introduces a new ordering for graphs that is multi-scale, coarse-to-fine. We define autoregressive learning for graphs as next-scale prediction instead of traditional next-node and next-edge approaches, thereby infusing autoregressive properties with graphs’ structural constraints. As exemplified in Figure 1, our method first encodes a graph into multiple scales of latent representations and learns to generate scales autoregressively, starting from a singular token map and progressively expanding its resolution. At each step, a transformer predicts the next-resolution token map, conditioned on all previous scales and a class label [C] for conditional synthesis.
MAG poses significant improvements over traditional AR methods on both generic and molecular graph datasets within all experiments. Notably, it closes the gap with diffusion baselines, surpassing them in data efficiency, training time, inference speed, and scalability, with only little overhead in performance. This provides strong evidence regarding the effectiveness of next-scale prediction for scalable and high-quality graph generation. In summary, our contributions are threefold:
@@
    A novel next-scale prediction framework for hierarchical graph generation that overcomes the limitations of traditional autoregressive approaches on graphs properties.
@@
    Experiments on diverse graph datasets, demonstrating significant improvements over AR baselines and little performance overhead compared to state-of-the-art diffusion methods.
@@ 2 Related Work
Diffusion models achieve permutation-invariance by parameterizing the denoising/score function with a permutation equivariant network, at the cost of up to thousands of sampling steps. EDP-GNN [6] is the first work to adapt score-based models to graph generation by representing graphs as matrices with continuous values. GDSS [7] generalizes this approach by leveraging SDE-based diffusion and replacing adjacency matrices with node and edge features. However, by using continuous-state diffusion, these approaches ignore graphs’ discrete nature, resulting in fully connected graphs. DiGress [8] is the first to apply discrete-state diffusion, achieving significant improvements over continuous methods; yet, they require additional structural and domain-specific features to break symmetries and achieve high generation quality, further increasing computational costs. In response, SwinGNN [11] argues that learning exchangeable probabilities with equivariant architectures is challenging, instead proposing fixed node orderings and non-equivariant score functions to alleviate complexity while maintaining competitive performance.
@@
Traditional AR models construct graphs by adding nodes and edges sequentially. Although this approach aligns with the discrete nature of graphs, it faces a fundamental challenge, as there is no inherent order in graph generation. Various strategies have been explored to address this by enforcing some node ordering and approximating the marginalization over permutations. Li et al. [16] suggest using random or deterministic empirical orderings. GraphRNN [3] aligns permutations with Breadth-First Search (BFS) ordering through a many-to-one mapping. GRANs [14] propose an approach to marginalize over a family of canonical node orderings, such as node degree descending, BFS trees rooted at the highest degree node, and k-core orderings. Chen et al. [17] eliminate ad-hoc orderings altogether by modeling the conditional probability of orderings with another AR model to estimate the marginalized probabilities for both the generative and ordering-selection models. Despite these efforts, all such methods remain fundamentally limited by the need to impose artificial node orderings.
@@
Hybrid models attempt to combine diffusion and AR models to balance performance and speed. PARD [10] decomposes graph generation into blocks before applying diffusion within each block and AR between them, while ensuring permutation invariance through partial ordering. However, block-wise diffusion may introduce implementation complexity and remain limited by sampling efficiency [18]. In contrast, GraphArm [9] employs an AR diffusion model where, at each forward step, exactly one node and its adjacent edges decay to absorbing states based on a random node ordering, making the model both permutation- and order-sensitive. Coarse-to-fine graph generation has also been explored. Bergmeister et al. [19] generate a full graph from a single node through localized denoising diffusion and spectral conditioning, achieving high performance but with generation times similar to those of full diffusion models. Meanwhile, JT-VAE [4] and HierVAE [20] generate graphs hierarchically using chemically valid molecular substructures, making these methods only viable on molecular tasks. Lastly, HiGGs [21] introduce a hierarchical framework that leverages GNNs with conditional generation capabilities to sample graphs across multiple resolution levels and extend the scale of the generated graphs with respect to training inputs.
@@ 3 Method
Consider a graph defined by the quadruple $G=(\mathcal{V},E,\mathbf{X},\mathbf{E})$, where $\mathcal{V}$ denotes the set of $N$ vertices, $E\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges, $\mathbf{X}\in\mathbb{R}^{N\times D}$ is the matrix containing $D$-dimensional node features, and $\mathbf{E}\in\mathbb{R}^{N\times N\times F}$ comprises edge attributes with feature dimensions $F$. Given a collection of $M$ observed graphs, $\mathcal{G}=\{G_{i}\}_{i=1}^{M}$, the task of graph generation is to model the underlying distribution $p(\mathcal{G})$ in order to generate new graphs $G_{\text{sample}}\sim p(\mathcal{G})$.
@@ 3.1 Preliminary: Autoregression by Scale
AR models represent the joint distribution over $N$ random variables by employing the chain rule of probability. In particular, they decompose the generation process into sequential steps, where each step determines the subsequent action based on the current subgraph. Traditionally, previous works have taken a next-token prediction approach, generating the graph one node and its incident edges at a time. Specifically, at step $i$, we introduce node $G_{i}^{\pi}$ along with any new edges connecting it to nodes in $\{G_{1}^{\pi},G_{2}^{\pi},\ldots,G_{i-1}^{\pi}\}$. The general formulation of AR models for graphs is given by:
@@
Alternatively, rather than employing next-token prediction, we suggest utilizing next-scale prediction, wherein the autoregressive unit is a latent representation of the entire graph rather than a single token. To implement this, we first obtain latent representations at multiple scales and train a transformer to generate these latent maps. Similarly to Tian et al. [1], this leads to a two-stage architecture: (1) a VQ-VAE tokenizer that quantizes the input graphs into latent maps at multiple scales (detailed in Section 3.2), and (2) a transformer that predicts next-scale latent maps (discussed in Section 3.3).
@@ 3.2 Permutation Equivariant Multi-Scale Tokenizer
We propose a permutation-equivariant quantized autoencoder to encode a graph $G$ to $K$ multi-scale discrete latent maps $R=\{r_{1},r_{2},...,r_{K}\}$, where each scale $r_{k}$ depends only on its prefix $\{r_{1},r_{2},...,r_{k-1}\}$. For quantization, a shared codebook $Z$ is used accross all scales, ensuring that each token in $r_{k}$ belongs to the same vocabulary. An overview of the tokenizer is given in Figure 2.
@@
where $C$ is the latent’s feature dimension defined empirically. In order to encode the graph’s node and edge features into a unified latent space $\mathbb{R}^{N\times C}$, we leverage a series of MPNNs, such that edge features are propagated within node features before compressing node dimensions to $C$ and dropping remaining edges.
@@
where $\text{Select}(Z,v)$ denotes selecting the $v^{th}$ vector in codebook $Z$, $k\leq K$ and $i\leq N$. $\mathcal{Q}(\cdot)$ is trained by fetching all $q^{(k,i)}$, given $F$, and minimizing the distance between $q$ and $f$.
@@ 3.3 Autoregressive Graph Generation Through Next-Scale Prediction
We define the autoregressive modeling on graphs by shifting from next-node and next-edge predictions to next-scale predictions. Here, the autoregressive unit is an entire token map, rather than a single token. Given the quantized scales $\{r_{1},r_{2},\ldots,r_{K}\}$, the autoregressive likelihood is formulated as:
@@
A level embedding is learned to encode the scale at which a token operates within the input structure. This is implemented similarily to GPT’s segment embedding [22], such that learned signals enhance the model’s capacity to capture multi-scale contextual dependencies.
@@
During training, we uniformly sample $K$ scale sizes to enhance the model’s generalization ability to perceive any scale. Specifically, we set the maximum number of scales to $K_{\text{max}}$ based on the task, and fix the first scale to $1$ and the last to $N$, where $N$ is the desired graph size. Intermediate scale sizes are uniformly sampled. During inference, we set all scales to predefined sizes (e.g., $R=\{1,2,4,6,9\}$) and generate them autoregressively, such that $\forall r_{k}\in R,r_{k}\leq N$ and the final scale $r_{K}=N$ (e.g., truncating to $\{1,2,4,5\}$ for a graph of $5$ nodes).
@@
We adopt a standard decoder-only transformer architecture with adaptive layer normalization (AdaLN). For class-conditional synthesis, the class embedding is used as the start token [C] and the condition of AdaLN. We do not use any advanced positional embeddings, such as those delineated in Graphormer [23] or DiGress [8].

### v2

# Diffusion-Free Graph Generation with Next-Scale Prediction

@@ Abstract
Autoregressive models excel in efficiency and plug directly into the transformer ecosystem, delivering robust generalization, predictable scalability, and seamless workflows such as fine-tuning and parallelized training. However, they require an explicit sequence order, which contradicts the unordered nature of graphs. In contrast, diffusion models maintain permutation invariance and enable one-shot generation but require up to thousands of denoising steps and additional features for expressivity, leading to high computational costs. Inspired by recent breakthroughs in image generation, especially the success of visual autoregressive methods, we propose MAG, a novel diffusion-free graph generation framework based on next-scale prediction. By leveraging a hierarchy of latent representations, the model progressively generates scales of the entire graph without the need for explicit node ordering. Experiments on both generic and molecular graph datasets demonstrated the potential of this method, achieving inference speedups of up to three orders of magnitude over state-of-the-art methods, while preserving high-quality generation.
Multi-scale, Hierarchical, Next-scale, Diffusion-free, Autoregressive, Graph, Molecules, Generation, Transformer, Generative AI, Machine Learning
@@ 1 Introduction
Despite significant progress in generative models for language and images, graph generation remains challenging due to its inherent combinatorial nature. Specifically, graphs are naturally high-dimensional and discrete with varying sizes, contrasting with the continuous and fixed-size techniques that cannot be directly applied to them. Furthermore, rich substructures in graphs necessitate an expressive model capable of capturing higher-order motifs and interactions.
Several graph generative models have been proposed to address some of these challenges, undertaking approaches based on autoregressive (AR) models (You et al., 2018), varational autoencoders (VAE) (Jin et al., 2018), generative adversarial networks (De Cao & Kipf, 2018), and diffusion models (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023). Among these, diffusion and autoregressive models excel in performance and offer complementary strengths—diffusion’s permutation invariance and AR’s efficiency—that make them especially promising avenues (Kong et al., 2023; Zhao et al., 2024; Yan et al., 2024).
On one hand, diffusion models offer the ability to achieve exchangeable probability in combination with permutation equivariant networks, under certain conditions (Niu et al., 2020; Jo et al., 2022). However, given the high-dimensional nature of graphs and their complex internal dependencies, directly modeling the joint distribution of all nodes and edges presents significant challenges. Recent work by Zhao et al. (2024) demonstrates that, as a cost for permutation invariance, capturing the full joint distribution and solving the transformation difficulty via diffusion requires thousands of sampling and denoising steps as well as additional node-, edge-, and graph-level features, such as eigenvectors, to break symmetries and achieve high generation quality, rendering diffusion a promising but very expensive approach.
On the other hand, in addition to their improved efficiency, studies into the success of AR models have highlighted their scalability and generalizabilty. The former, as exemplified by scaling laws (Henighan et al., 2020), allows the prediction of large models’ performance to better guide resource allocation during training, while the latter, as evidenced by zero-shot and few-shot learning (Sanh et al., 2021), underscores the adaptability of unsupervised-trained models to unseen tasks. Furthermore, the aforementioned discrete and varying-size properties of graphs suggest a promising fit for autoregressive modeling. However, unlike natural language sentences with an inherent left-to-right ordering, the order of nodes in a graph must be explicitly defined for unidirectional autoregressive learning. Previous AR methods (You et al., 2018; Liao et al., 2020; Shi et al., 2020) attempted to enforce node ordering or approximate the marginalization over permutations such that, once flattened, one can train an autoregressive model to maximize its likelihood via next-token prediction. This gives rise to the three main issues of autoregressive graph generation:
@@
    Long-range dependency. Generating nodes one at a time results in sequential dependencies that make it difficult to capture long-range interactions and global structures. Furthermore, early predictions errors can propagate throughout the generation process, and dependencies may decay when working on larger graphs.
@@
    Forced node ordering. Enforcing a specific node ordering is at odds with the inherent invariance property of graphs, forcing the model to learn over a specific ordering, thereby increasing complexity and reducing generalization.
@@
    Inefficiency. Generating a graph sequence $(x_{1},x_{2},...,x_{N})$ with a conventional self-attention transformer incurs $\mathcal{O}(n)$ autoregressive steps and $\mathcal{O}(n^{3})$ computational cost (see Appendix B).
Drawing inspiration from recent breakthroughs in image generation—notably the Visual Autoregressive Modeling (VAR) framework (Tian et al., 2024; Ren et al., 2025)—our proposed method suggests a new ordering for graphs that is multi-scale, coarse-to-fine. We introduce MAG, a novel transformer-based graph generation framework based on next-scale prediction instead of traditional next-node and next-edge approaches, thereby infusing autoregressive properties with graphs’ structural constraints. As shown in Figure 1, our method first encodes a graph into multiple scales of latent representations and learns to generate scales autoregressively, starting from a singular token and progressively expanding its resolution. At each step, a transformer predicts the next-resolution token map, conditioned on previous scales and a class label [C] for conditional synthesis.
MAG poses significant improvements over traditional autoregressive methods on both generic and molecular graph datasets. Notably, it reduces the performance gap with diffusion baselines, surpassing them in wall-clock training and inference time, with strong performance. This provides strong evidence regarding the effectiveness of next-scale prediction for high-quality graph generation. In summary, our contributions are threefold:
@@
    A novel next-scale prediction framework for hierarchical graph generation that overcomes the limitations of traditional AR approaches on graph properties.
@@
    Experiments on diverse graph datasets, demonstrating significant improvements over AR baselines and promising performance when compared to state-of-the-art diffusion methods.
@@ 2 Related Work
Diffusion models achieve permutation-invariance by parameterizing the denoising/score function with a permutation equivariant network, at the cost of up to thousands of sampling steps. EDP-GNN (Niu et al., 2020) is the first work to adapt score-based models to graph generation by representing graphs as matrices with continuous values. GDSS (Jo et al., 2022) generalizes this approach by leveraging SDE-based diffusion and replacing adjacency matrices with node and edge features. However, by using continuous-state diffusion, these approaches ignore graphs’ discrete nature, resulting in fully connected graphs. DiGress (Vignac et al., 2023) is the first to apply discrete-state diffusion, achieving significant improvements over continuous methods. Yet, they require additional structural and domain-specific features to break symmetries and achieve high generation quality, further increasing computational costs. In response, SwinGNN (Yan et al., 2024) argues that learning exchangeable probabilities with equivariant architectures is challenging, instead proposing fixed node orderings and non-equivariant score functions to alleviate complexity while maintaining competitive performance.
@@
Traditional AR models construct graphs by adding nodes and edges sequentially. Although this approach aligns with the discrete nature of graphs, it faces a fundamental challenge that there is no inherent order in graph generation. Various strategies have been explored to address this by enforcing some node ordering and approximating the marginalization over permutations. Li et al. (2018) suggest using random or deterministic empirical orderings. GraphRNN (You et al., 2018) aligns permutations with Breadth-First Search (BFS) ordering through a many-to-one mapping. GRANs (Liao et al., 2020) propose an approach to marginalize over a family of canonical node orderings, such as node degree descending, BFS trees rooted at the highest degree node, and k-core orderings. Chen et al. (2021) eliminate ad-hoc orderings altogether by modeling the conditional probability of orderings with another AR model to estimate the marginalized probabilities for both the generative and ordering-selection models. Despite these efforts, all such methods remain fundamentally limited by the need to impose artificial node orderings.
@@
Several recent works seek to blend the strengths of diffusion and autoregressive (AR) paradigms. PARD (Zhao et al., 2024) first partitions the graph into blocks, applies diffusion within each block and AR between them, and enforces permutation invariance via partial ordering. While effective, its block-wise diffusion introduces extra implementation complexity and still incurs substantial sampling costs (Hou et al., 2024). GraphArm (Kong et al., 2023) instead embeds diffusion steps into an AR framework—at each forward pass, a single node and its edges stochastically decay—making the model both permutation- and order-sensitive. Beyond these hybrids, pure coarse-to-fine strategies have also been explored: Bergmeister et al. (Bergmeister et al., 2024) grow a full graph from one seed node using localized denoising diffusion and spectral cues, matching diffusion-only runtimes but retaining their high sampling overhead. In the molecular domain, JT-VAE (Jin et al., 2018) and HierVAE (Jin et al., 2020) build molecules by assembling chemically valid substructures, and HiGGs (Davies et al., 2023) hierarchically samples entire graphs across multiple resolutions. However, all of these rely on either block-level diffusion, node-level decay processes, or strong domain priors (e.g., motif libraries or spectral conditioning). In contrast, our framework is fully domain-agnostic and permutation-equivariant, leveraging multi-scale quantized latent maps with a generic transformer—without any hand-crafted substructure priors.
@@ 3 Method
Consider a graph defined by the quadruple $G=(\mathcal{V},E,\mathbf{X},\mathbf{B})$, where $\mathcal{V}$ denotes the set of $N$ vertices, $E\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges, $\mathbf{X}\in\mathbb{R}^{N\times D}$ is a matrix of $D$-dimensional node features, and $\mathbf{B}\in\mathbb{R}^{N\times N\times F}$ comprises edge attributes with dimensions $F$. Given a collection of $M$ observed graphs, $\mathcal{G}=\{G_{i}\}_{i=1}^{M}$, the task of graph generation is to model the underlying distribution $p(\mathcal{G})$ in order to generate new graphs, $G_{\text{sample}}\sim p(\mathcal{G})$.
@@ 3.1 Preliminary: Autoregression by Scale
AR models represent the joint distribution over $N$ random variables by employing the chain rule of probability. In particular, they decompose the generation process into sequential steps, where each step determines the subsequent action based on the current subgraph. Traditionally, previous works have taken a next-token prediction approach, generating the graph one node and its edges at a time. Specifically, at step $i$, we introduce node $G_{i}^{\pi}$ along with any new edges connecting it to nodes in $\{G_{1}^{\pi},G_{2}^{\pi},\ldots,G_{i-1}^{\pi}\}$. The general formulation of AR models for graphs is given by:
@@
Alternatively, rather than employing next-token prediction, we propose utilizing next-scale prediction, wherein the autoregressive unit is a latent representation of the entire graph rather than a single token. To implement this, we first obtain latent representations at multiple scales and train a transformer to generate these latent maps. Similarly to Tian et al. (2024), this leads to a two-stage architecture: (1) a VQ-VAE tokenizer that quantizes the input graphs into latent maps at multiple scales (refer to Section 3.2), and (2) a transformer that predicts next-scale latent maps (refer to Section 3.3).
@@ 3.2 Permutation Equivariant Multi-Scale Tokenizer
We propose a permutation-equivariant quantized autoencoder to encode a graph $G$ to $K$ multi-scale discrete latent maps $R=\{r_{1},r_{2},...,r_{K}\}$, where each scale $r_{k}$ depends only on its prefix $\{r_{1},r_{2},...,r_{k-1}\}$. For quantization, a shared codebook $Z$ is used across all scales, ensuring that each token in $r_{k}$ belongs to the same vocabulary. An overview of the tokenizer is given in Figure 2.
@@
where $C$ is the latent’s feature dimension defined empirically. In order to encode the graph’s node and edge features into a unified latent space $\mathbb{R}^{N\times C}$, we leverage a series of standard MPNNs, such that edge features are propagated within node features before compressing node dimensions to $C$ and dropping remaining edges.
Concretely, each MPNN layer performs message passing by first computing learned linear transformations of neighbor node embeddings and their connecting edge attributes, aggregating these via a sum operation, and then applying a ReLU nonlinearity. We stack $L_{e}$ such layers—with residual connections and layer normalization—to ensure stable training and encode higher-order structural patterns throughout the graph within nodes.
@@
where $\text{Select}(Z,v)$ denotes selecting the $v^{th}$ vector in codebook $Z$, and $i\leq N$. Specifically, $\mathcal{Q}(\cdot)$ is trained by fetching all $q^{(k,i)}$ and minimizing the distance between $q$ and $f$.
@@
We implement $\mathcal{D}(\cdot)$ with $L_{d}$ standard graph convolutional layers, interleaved with ReLU activations and layer normalization to stabilize training and capture higher-order structures. After the final GCN layer, edge logits are recovered by passing all concatenated pairs of node embeddings through an MLP layer, and node features are mapped back to input dimensions, $\mathbb{R}^{D}$, via a linear projection.
@@ 3.3 Autoregressive Graph Generation Through Next-Scale Prediction
We define the autoregressive modeling on graphs by shifting from next-node and next-edge predictions to next-scale predictions. Here, the autoregressive unit is an entire token map, rather than a single token. Given the quantized scales $\{r_{1},\ldots,r_{K}\}$, the autoregressive likelihood is expressed as:
@@
A level embedding is learned to encode the scale at which a token operates within the input structure. This is implemented similarily to GPT’s segment embedding (Devlin et al., 2019), such that learned signals enhance the model’s capacity to capture multi-scale contextual dependencies. However, it is important to note that no additional positional encoding is used within scales, in order to leverage Attention’s permutation-invariance property.
@@
During training and inference, we select scale sizes from a predefined set of lengths. Specifically, we fix the first scale to $1$ and the final scale to $N$, where $N$ is the target graph size. Intermediate scales are chosen from the predefined set (e.g., $R=\{1,2,4,6,9\}$), such that $\forall r_{k}\in R,r_{k}\leq N$; and the final scale $r_{K}=N$ (e.g., truncating to $\{1,2,4,5\}$ for a graph of $5$ nodes). Typically, we set the number of scales to be logarithmic in the number of nodes. Finally, to handle varying graph sizes and enable parallel generation, padding is used with appropriate masking.
@@
We adopt a standard decoder-only Transformer architecture with adaptive layer normalization (AdaLN), and input tokens are embedded via a linear layer. Unlike Graphormer (Ying et al., 2021) or DiGress (Vignac et al., 2023), we do not use any additional positional embeddings. For class-conditional synthesis, a class embedding serves both as the start token [C] and the conditioning input for AdaLN. Then, at each autoregressive step, the next scale is predicted such that all tokens in that scale are generated in parallel through a series of multi-head self-attention blocks. Each output embedding is passed through a classifier head to predict a categorical distribution over the quantization codebook, from which the most probable codeword is selected using a softmax operation. Dropout and layer normalization are employed throughout the architecture to aid convergence. Additionally, query and key vectors are normalized to unit vectors prior to computing attention scores to improve training stability. To support reproducibility, the hyperparameters used for training and inference are listed in Appendix A.
Sampling at inference
At inference, we apply top-$k$ and top-$p$ (nucleus) sampling to introduce diversity and control over the generation process. These methods are applied to the predicted categorical distributions at each decoding step.
Figure 4: Samples of generated generic and molecular graphs by MAG.